OpenStudy (anonymous):

A large emerald with a mass of 982.7 grams was recently discovered in a mine. If the density of the emerald is 2.76grams over centimeters cubed, what is the volume? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point [Fill in the Blank]

OpenStudy (anonymous):

@Jack1 can u help me when ever u get the chance (:

OpenStudy (jack1):

sure mass = 982.7 grams Density = 2.76 g/cm^3 Volume = ? so Density = mass per unit of volume \(\Large D = \frac mV \) if m = 982.7 and D = 2.76 V = ??? \(\Large D = \frac mV \) \(\Large 2.76 = \frac {982.7}V \) Can you solve for V from here?

OpenStudy (anonymous):

sorry i was from my computer again but would it be 276.92??

OpenStudy (jack1):

no i got a different answer, how did u calculate that?

OpenStudy (anonymous):

sorry i think i calculated it wrong

OpenStudy (jack1):

s'ok \(\Large 2.76 = \frac {982.7}V\) multiply both sides by V \(\Large 2.76 \times V = \frac {982.7}V \times V\) V's now cancel out on the Right Hand Side, so you left with \(\Large 2.76 \times V = 982.7\) now divide both sides by 2.76 \(\Large \frac {2.76 \times V}{2.76} = \frac {982.7}{2.76}\) 2.76 over 2.76 on Left Hand Side = 1 \(\Large V = \frac {982.7}{2.76}\) so answer = ?

OpenStudy (anonymous):

i divided and got 356.05 was i suppose to divided ??

OpenStudy (jack1):

356.05 = perfect answer, nicely done

OpenStudy (geekfromthefutur):

Giid job Moo Moo

OpenStudy (geekfromthefutur):

Good

OpenStudy (anonymous):

Thank you ^w^ and really geek lol

OpenStudy (geekfromthefutur):

hurry up amanda

OpenStudy (jack1):

nice spelling there geek ;P

OpenStudy (anonymous):

hey jack would it be ok if u could help me with 4 more questions :D

OpenStudy (geekfromthefutur):

Ty Goku Jack

OpenStudy (jack1):

lol ;)

OpenStudy (geekfromthefutur):

ill help too

OpenStudy (jack1):

yo, sure, close this one off tho, it's all LaTex'd up again

OpenStudy (geekfromthefutur):

I hate the latex subject

OpenStudy (anonymous):

oh ok ^w^ thank you